Shift summary - Eve 12/14/99

Julie Whitmore (jaws@fnal.gov)
Wed, 15 Dec 1999 00:02:41 -0600 (CST)

Shift Summary for eve shift, 14/Dec/99

On Shift:Herman, Julie

(Julie decides to take 799 rather than 832 data during this shift.)
Hours of beam available at all intensities: 8.2hrs
Hours of beam available at nominal intensity: 8.0hrs
Hours of E799 data (special runs not included): 6.8hrs
Hours of special runs (laser scans, muon runs...): 0.0hr

Accounting of lost time:
Trigger latch problems. Stream 5 permit problems.
Beamline magnet trips. Rad trips.

========================================================================
Runs ended this shift:
None

Tapes brought to FCC this shift:
(including tapes for online split)
None

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Steady beam. Rough data taking because of frontend problems. See
Logbook XXIII pp.72-74 for details.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
16:15 Onsplit request - UPF057 on ktev2

15219 51-53 16:25 YTF DC1-6 up is hot. Page JasonL.
Banana problem. Fixed with reset.

67-73 16:45 Low intensity (6E12). Herman phones
MCR. They are doing a MI studies.

180 19:15 SDAQ mon_caen process is dead. SashaL tries
to restart. TRDs are holding HV but cannot
be monitored. Flying blind. Caen supplies
need to power cycled. Valerie put on call
list.

202-241 19:45 Lots of ktevana errors which cannot be
fixed by kfrending. SashaL tracks down
problem to Crate 8, the trigger mask crate.
RickK paged. He suggests we bypass the
crate. Done.
MANY MONITORING HISTOGRAMS WILL BE BAD.

20:59 Onsplit request - UPK030 on ktev1.

265-278 21:12 Lost stream 5 permit. Kfrending & cycling
laser repeater crate doesn't seem to clear
it. Page SashaL who advises me to continue
kfrending. Eventually it clears.

22:30 RickT phones to report MU2 ADC channels
33 and 34 are dead. He'd like to swap the
modules during long down time tomorrow.

========================================================================
Follow up Notes on Previous Shift
Accidental 90 monitoring cable (#4 --> #1) switch. Page BobT. Not done.

========================================================================
Items to Follow up on Next Shift

Call list:
Page BobT if beam is down for >= 1hr on day or eve shifts.
Page Sasha Gol. and JasonL if beam is down for > 30min on day or eve shifts.
Page SashaL and Valerie (x3334) if beam is down for > 15min on any shift.

Trigger monitoring plots are screwed up because crate 8 is out of readout.
MU2 channels 33/34 missing in plot. Generates frowny.
Run 15219 - YTF DC1 UP - 6 is hot
hccmon plot is bad; aftermath of readout problems.

TRD HV Caen monitoring is not running. Need to monitor TRDs by looking at
TRD hit plot on detector monitor screen. See that hit count does not drift
down to 0.